# Knots and variance ordering of sequential Monte Carlo algorithms

### v1

@@ 6.4 Optimality of adapted terminal knots
Example 6.16 (A sequence of adapted knot-models, continued).
6.5 Estimating normalising constants
One of the most useful cases for terminal knots is when the normalising constant of the Feynman–Kac model, $\hat{\gamma}_{n}(1)$, of the model is of primary interest. If $\varphi=1$ is the only test function of interest then it is possible to specify a simplified Feynman–Kac model, which we detail in Example 6.17. Note that we specify a knotset which includes a terminal knot in this example, which we call a terminal knotset.
Consider a $\phi$-extended model $\mathcal{M}^{\phi}\in\mathscr{M}_{n}$ and knotset $\mathcal{K}=(R_{0:n},K_{0:n})$ where $\phi=1$. The knot-model $\mathcal{M}^{\ast}=\mathcal{K}\ast\mathcal{M}^{\phi}=(M^{\ast}_{0:n},G^{\ast}_{0:n})$ satisfies $M^{\ast}_{0}=R_{0}$, $M^{\ast}_{p}=K_{p-1}^{G_{p-1}}R_{p}$ for $p\in[1\,{:}\,n-1]$, $M_{n}^{\ast}=K_{n-1}^{G_{n-1}}R_{n}\otimes K_{n}^{G_{n}}$, $G^{\ast}_{p}=K_{p}(G_{p})$ for $p\in[0\,{:}\,n-1]$, and $G_{n}^{\ast}=K_{n}(G_{n})\otimes 1$.
The model $\mathcal{M}^{\dagger}=(M^{\dagger}_{0:n},G^{\dagger}_{0:n})$ with
$M_{0}^{\dagger}$ $=R_{0},$ $\quad G_{0}^{\dagger}$ $=K_{0}(G_{0}),$
Note that $K_{n}^{G_{n}}$ does not need to be simulated in the model $\mathcal{M}^{\dagger}$. The asymptotic variance for model $\mathcal{M}^{\dagger}$ can also be further reduced by apply more knots. A special case of Example 6.17 is using the terminal adapted knotset. In this case, $M_{0}^{\dagger}=\delta_{0}$ and so long as $G_{0}^{\dagger}=M_{0}(G_{0})$ is accounted for elsewhere in the algorithm, the first iteration of the SMC algorithm does not need to be run.
7 Applications and examples
7.1 Particle filters with ‘full’ adaptation
A particle filter with ‘full’ adaptation adapts each Markov kernel in the Feynman–Kac model to the current potential information through twisting. Originally proposed as a type of auxiliary particle filter by Pitt and Shephard (1999), its modern interpretation does away with auxiliary variables though it is still often referred to as a fully-adapted auxiliary particle filter. It is popular due to its empirical performance and its derivation which is motivated by identifying locally (i.e. conditional) optimal proposal distributions at each time step. We refer to this algorithm as a particle filter with ‘full’ adaptation. The Feynman–Kac model for such an algorithm is described in Example 7.1.
Example 7.1 (Particle filter with ‘full’ adaptation).
$M_{n}^{\mathrm{F}}$ $=M_{n}^{G_{n}},$ $\quad G_{n}^{\mathrm{F}}$ $=1.$
The adapted knot-model in Example 3.11 and the particle filter with ‘full’ adaptation in Example 7.1 share the same constituent twisted Markov kernels $M_{p}^{G_{p}}$ and expected potential functions $M_{p}(G_{p})$, but differ in where these elements are located in time. A further crucial difference is that adapted knot-model is not adapted to the terminal potential. Our theory on knots has shown that adapted knot-models order the asymptotic variance for all relevant test functions, whilst Johansen and Doucet (2008) contained a counter example to such a result for the particle filter with ‘full’ adaptation, which they referred to as ‘perfect’ adaptation. We restate the model for the counter example in Example 7.2, and will demonstrate how adapted knot-models guarantee an asymptotic variance reduction where the fully-adapted particle filter does not.
Example 7.2 (Binary model of Johansen and Doucet, 2008).
Let $\mathcal{B}=(M_{0:1},G_{0:1})$ be a Feynman–Kac model with
[…]
respectively, and potential functions
$G_{t}(x_{t})=\begin{cases}1-\varepsilon&\text{if}~x_{t}=y_{t},\\ \varepsilon&\text{if}~x_{t}=1-y_{t},\\ \end{cases},\quad t\in\{0,1\}$
with fixed observations $y_{t}$ such that $y_{0}=0$ and $y_{1}=1$.
Figure 2 compares the asymptotic variances of the bootstrap PF, PF with ‘full’ adaptation, and adapted knotset PF in Example 7.2 with $\varepsilon=0.25$ and $\delta\in(0,1)$ both analytically and empirically. The Figure considers the particle approximation $\hat{\eta}_{1}^{N}(\varphi)$ with $\varphi(x)=x$, and replicates Figure 2 in Johansen and Doucet (2008) with the addition of the adapted knotset PF. The adapted knotset PF has underlying model with $M_{0}^{\ast}=\delta_{0}$, $M_{1}^{\ast}=M_{0}^{G_{0}}M_{1}$, $G_{0}^{\ast}=M_{0}(G_{0})$, and $G_{1}^{\ast}=G_{1}$. We can see that the adapted knotset PF outperforms the other PFs in this regime, whilst the bootstrap PF almost always has lower variance that the PF with ‘full’ adaptation when $\varepsilon=0.25$. The existence of regimes where the bootstrap PF is better than the PF with ‘full’ adaptation constitutes the counter example of Johansen and Doucet (2008).
Figure 3 compares the analytical asymptotic variances of the PF with ‘full’ adaptation and adapted knotset PF relative to the bootstrap PF for $\varepsilon\in\{0.05,0.1,0.2,0.4,0.5\}$ (and symmetrical results) and $\delta\in(0,1)$. We see that the adapted knotset PF is always less than or equal to zero, demonstrating the dominance over the bootstrap PF, whilst the PF with ‘full’ adaptation can be better or worse than the bootstrap depending on the regime. We also note that the PF with ‘full’ adaptation can outperform the adapted knotset PF for some parameter values. However, nothing can be said in general as this is specifically for the test function $\varphi(x_{1})=x_{1}$. We have only shown knot-models or equivalent specifications have guaranteed variance ordering for all relevant test functions.
Figure 3: Analytical ($N\rightarrow\infty$) asymptotic variance of filter with ‘full’ adaptation and adapted knotset particle filter relative to the bootstrap particle filter for $\hat{\eta}_{1}^{N}(\varphi)$ in Example 7.2 with $\varepsilon=0.25$ and $\varphi(x)=x$. The empirical variance is estimated from 50,000 independent replications. Negative values indicate an improvement over the bootstrap particle filter.
We also compare normalising constant estimation in Example 7.2 for the Bootstrap PF, PF with ‘full’ adaption, and PF with terminal knotset. For the latter particle filter we use the simplified version in Example 6.17 with adapted knots.
Figure 4: Analytical ($N\rightarrow\infty$) and empirical (N=100, [^0] N=100{,}000) asymptotic variance of bootstrap particle filter, filter with ‘full’ adaptation, and adapted knotset particle filter for $\hat{\gamma}_{1}^{N}(1)$ in Example 7.2 with $\varepsilon=0.25$. The empirical variance is estimated from 50,000 independent replications.
Figure 4 compares the asymptotic variances of each particle filter for Example 7.2 with $\varepsilon=0.25$ and $\delta\in(0,1)$ both analytically and empirically. The Figure illustrates the particle approximation of the normalising constant, $\hat{\gamma}_{1}^{N}(1)$. The terminal adapted knotset PF has underlying model with $M_{0}^{\ast}=\delta_{0}$, $M_{1}^{\ast}=M_{0}^{G_{0}}$, $G_{0}^{\ast}=M_{0}(G_{0})$, and $G_{1}^{\ast}=M_{1}(G_{1})$ as in Example 6.17. When estimating the normalising constant, the terminal adapted knotset PF guarantees an asymptotic variance reduction, whilst the PF with ‘full’ adaptation does not. In this sense, Figure 4 serves as a counter-example to the notion that the PF with ‘full’ adaptation guarantees variance reduction for normalising constant estimation. Both models use identical constituent elements and result in similar algorithmic implementations. As such, we might expect the asymptotic variance of these PFs to be equal across parameter values. In fact, this is not the case, and only the terminal adapted knotset PF guarantees a variance reduction.
Overall, our approach to variance reduction with knots explains why the particle filter with ‘full’ adaptation has good empirical performance in many different contexts: It is only slightly different to a model (Example 3.11) that we can guarantee an asymptotic variance ordering for all relevant functions. Thus, we provide a cohesive explanation for the counter example in Johansen and Doucet (2008) by clarifying that it is the adaptation to the terminal time and placement of the constituent elements of the model that restricts the variance reduction guarantee, not the use of adaptation that is only locally (i.e. conditionally) optimal. Further, we have demonstrated how a minor modification to the particle filter with ‘full’ adaptation, Example 3.11, guarantees variance ordering for all relevant test functions which has practical significance.
7.2 Marginalisation as a knot
Model marginalisation in SMC is an well-known variance reduction technique that can be viewed as a special case of knots. Often referred to as Rao–Blackwellisation, the procedure involves analytically marginalising part of the state-space of the Feynman–Kac model, thus reducing the dimensionality of the estimation problem. Historically, Rao–Blackwellisation has been applied to models where it is applicable at all time points, and justified in the case of sequential importance samplers by appealing to a reduction in the variance of the weights (Doucet, 1998; Doucet et al., 2000). However, this justification does not relate Rao–Blackwellisation to the variance of particle approximations from a modern SMC algorithm which, in comparison, uses resampling. Framing model marginalisation as a knot operation, we prove this procedure will order the asymptotic variance of SMC algorithms for all relevant test functions when the knot is applied to a non-terminal time point.
Example 7.3 (Marginalisation knot).
is a marginalisation knot and $M_{t}=RK$.
The kernels $U$ and $V$ partition the state-space $M_{t}$ is defined on. In particular, $U$ is the marginal distribution of the first component of the partition and $V$ is the conditional distribution of the second component of the partition. The result of apply a marginalisation knot to a model is considered next.
Example 7.4 (Model with marginalistion knot).
It is also instructive to note that a knot itself can be seen as model marginalisation. If $M_{t}=RK$ we could extend the Feynman–Kac model from $\mathcal{M}=(M_{0:n},G_{0:n})$ to $\mathcal{M}^{\prime}=(M_{0:n+1}^{\prime},G_{0:n+1}^{\prime})$ where $M_{t}^{\prime}=R$, $G_{t}^{\prime}=1$, $M_{t+1}^{\prime}=K$, $G_{t+1}^{\prime}=G_{t}$, and $M_{p+1}^{\prime}=M_{p}$ with $G_{p+1}^{\prime}=G_{p}$ for $p\in[t+1\,{:}\,n]$. Marginalising the state $X_{t+1}^{\prime}\sim K(x_{t}^{\prime},\cdot)$ in $\mathcal{M}^{\prime}$, and collecting the remaining terms, results in the model $\mathcal{K}\ast\mathcal{M}$ where $\mathcal{K}=(t,R,K)$. As such, a knot is marginalisation, or model extension followed by marginalisation, and our procedures and theory presents the most general case of this, as well as nuance around the use of knots at the terminal time.
7.3 Non-linear Student distribution state-space models
In this section we provide an numerical example to demonstrate the use of knots in practice and illustrate the connection between adapted knots and marginalisation knots. We consider a non-linear state-space model with latent variable driven by additive Student noise. The model uses non-linear functions $f_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ such that the latent space evolution can be described as
We will use the fact that a multivariate Student distribution can be represented as scale mixture of multivariate Gaussians with transformed $\chi^{2}_{\nu}$ distribution, that is a Chi-squared with $\nu$ degrees of freedom. Noting this construction, the Feynman–Kac form of this state-space model has Markov kernels satisfying
$M_{0}$ $=(\delta_{\mu}\otimes S)K$ (9)
To test the variance reductions possible for this model and knotset defined by $\mathcal{K}_{0},\ldots,\mathcal{K}_{n}$, we compare the bootstrap particle filter and the terminal knotset normalising constant particle filter in Example 6.17 over $R=200$ repetitions. Each particle filter used $N=2^{10}$ particles and adaptive resampling with threshold $\kappa=0.5$. We test the particle filters for five independent datasets simulated by the data generating process that varying by dimension $d\in[1\,{:}\,5]$. We fix the time horizon $n=10$, initial mean $\mu=0_{d}$, degrees of freedom $\nu=4$ and variance matrices $\Sigma=\Sigma^{\prime}=I_{d}$. We use a univariate non-linear function $g_{p}(x)=\frac{x}{2}+\frac{25x}{1+x^{2}}+8\cos\left(1.2p\right)$ (see Kitagawa, 1996, and references therein) to construct a multivariate non-linear function $f_{p}(x)=A[g_{p}(x_{1})\cdots g_{p}(x_{d})]^{\top}$ with matrix $A\in\mathbb{R}^{d\times d}$ having unit diagonal, off-diagonals elements set to half (when $d\geq 2$), and all other elements set to zero (when $d\geq 3$).
Figure 5: Box plots of (shifted) log normalising constant estimates of Bootstrap Particle Filter (BPF) and Terminal Knotset Particle Filter (TKPF). The estimates were shifted by a constant (for each $d$) for ease of comparison.
This example demonstrates a variance reduction for a class of models that appears to be unconsidered in the literature. From the view of adaptation, (Pitt and Shephard, 1999) showed that ‘full’ adaptation could be implemented with non-linear functions and additive noise. Whilst in this example, we condition on the $\chi^{2}_{\nu}$ distributed state and adapt only the Gaussian component in the extended state-space. Whereas from the marginalisation view, the general framework in Schön et al. (2005) does not include the possibility of marginalising a non-linear component, as we do here.
Many further generalisations of this example are possible. For example, for a model in the form of (9), a $\mathcal{K}_{p}$ knot leads to a tractable particle filter any conjugate $K$ and $G_{p}$ for arbitrary $S$. Such an $S$ could represent a scale mixture component, as in this example, further parts of the state-space, and need not be independent of past states.
8 Discussion
In terms of particle filter design, we have re-emphasised the importance of ‘full’ adaptation (or adapted knots) by explaining how the pitfall in the counter-example in Johansen and Doucet (2008) can be avoided by not adapting at the terminal time point. Further, given the guaranteed asymptotic variance ordering from knots, we recommend that every Feynman–Kac model be assessed to see if there are one or more tractable knots that can be applied. In such an assessment, the cost of computing $K(G_{p})$ and simulating from $K^{G_{p}}$ should also be considered.
There are several future research directions to explore. Adapted knots have a connection with twisted Feynman–Kac models (Guarniero et al., 2017; Heng et al., 2020). On an extended state-space, a knot can be thought of as decomposing a kernel of a Feynman–Kac model into $R_{t}$ and $K_{t}$, adapting $K_{t}$ to the potential $G_{t}$, and simplifying the resulting model. Whilst a “twist” at time $t$ decomposes the potential function into $\psi_{t}$ and $\frac{G_{t}}{\psi}$ and adapts $M_{t}$ to $\psi_{t}$. When $K_{t}=M_{t}$ and $\psi_{t}=G_{t}$ the knot-model and twist-model are equivalent up to a time-shift for $t<n$. Further developing this connection may suggest new methods for twisted Feynman–Kac models in particle filters. Similarly, look-ahead particle filters should also be considered (Lin et al., 2013). For now we note that, except for normalising constant estimation, it may be beneficial for a twisted model use $\psi_{n}=1$ as we know that terminal knots do not guarantee an asymptotic variance reduction for all relevant test functions.
The application of knots and related variance reductions for SMC samplers is also an area for future research. Consider an SMC sampler with target distribution $\hat{\eta}_{t-1}$ at time $t$ with $M_{t}=RK$. Further, let $R$ be a $k$-cycle of some $\hat{\eta}_{t-1}$-invariant MCMC kernel $P$, that is $R=P\otimes\cdots\otimes P$, and let $K$ be a random uniform selection over the path generated by $R$. Applying the knot $(t,R,K)$ to the resulting Feynman–Kac model recovers the proposed kernel and potential function used in the recent waste-free SMC algorithm (Dau and Chopin, 2022) at time $t$. Related connections to recent work on Hamiltonian Monte Carlo integrator snippets in an SMC-like algorithm (Andrieu et al., 2024) are also of interest.
Appendix A Supporting results
Proposition A.1 (Kernel untwisting).
@@ Appendix A Supporting results
$K(H)\cdot K^{H}(\varphi)=K(H\cdot\varphi).$
@@
Second case. If $x\in\mathsf{X}_{0}$ make three notes. Firstly, $K(H)(x)K^{H}(\varphi)(x)=K(H)(x)Q(\varphi)(x)$ by definition. Secondly, $H=0$ almost surely w.r.t. $K(x,\cdot)$ since $H$ is non-negative. Hence, $K(H\cdot\varphi)(x)=0$ using the standard measure-theoretic convention $0\times\infty=0$ (see for example, Billingsley, 1995, p. 199). Further, $K(H)(x)Q(\varphi)(x)=0$ for $Q(\varphi)(x)\in[-\infty,\infty]$ by the same convention. Therefore, $K(H)(x)K^{H}(\varphi)(x)=K(H\cdot\varphi)(x)=0$ for $x\in\mathsf{X}_{0}$. ∎
Proposition A.2 (Form of $Q_{p,n}$ with knots).
Suppose $\mathcal{M}^{\ast}=\mathcal{K}\ast\mathcal{M}$. If $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$ and $\varphi:\mathsf{X}_{n}\rightarrow[-\infty,\infty]$ then $Q_{p,n}^{\ast}(\varphi)=K_{p}Q_{p,n}(\varphi)$ almost everywhere for $p\in[0\,{:}\,n-1]$ and $Q_{n,n}^{\ast}(\varphi)=Q_{n,n}(\varphi)$.
@@
$Q_{p}^{\ast}(\varphi_{p})(y_{p-1})$ $=K_{p-1}(G_{p-1})(y_{p-1})K_{p-1}^{G_{p-1}}R_{p}(\varphi_{p})(y_{p-1})$ (10)
$=\int K_{p-1}(y_{p-1},\mathrm{d}x_{p-1})G_{p-1}(x_{p-1})R_{p}(\varphi_{p})(x_{p-1})$
for $y_{p-1}\in\mathsf{Y}_{p-1}$ by Proposition A.1. As for $Q_{n}^{\ast}$ we have
$Q_{n}^{\ast}(\varphi)(y_{n-1})$ $=K_{n-1}(G_{n-1})(y_{n-1})K_{n-1}^{G_{n-1}}M_{n}(\varphi)(y_{n-1})$
$=\int K_{n-1}(y_{n-1},\mathrm{d}x_{n-1})G_{n-1}(x_{n-1})M_{n}(\varphi)(x_{n-1})$
$=K_{n-1}Q_{n}(\varphi)(y_{n-1})$
for $y_{n-1}\in\mathsf{Y}_{n-1}$ by Proposition A.1. Now for the $Q_{p,n}^{\ast}$ terms, for p=n1p=n-1
@@
Assume $Q_{p+1,n}^{\ast}(\varphi)=K_{p+1}Q_{p+1,n}(\varphi)$ for a given $p\in[0\,{:}\,n-2]$. Then by (10), for $y_{p}\in\mathsf{Y}_{p}$,
$Q_{p,n}^{\ast}(\varphi)(y_{p})$ $=Q_{p+1}^{\ast}Q_{p+1,n}^{\ast}(\varphi)(y_{p})$
$=\int K_{p}(y_{p},\mathrm{d}x_{p})G_{p}(x_{p})R_{p+1}(x_{p},\mathrm{d}y_{p+1})K_{p+1}(y_{p+1},\mathrm{d}x_{p+1})Q_{p+1,n}(\varphi)(x_{p+1})$
$=\int K_{p}(y_{p},\mathrm{d}x_{p})G_{p}(x_{p})M_{p+1}(x_{p},\mathrm{d}x_{p+1})Q_{p+1,n}(\varphi)(x_{p+1})$
$=K_{p}Q_{p+1}Q_{p+1,n}(\varphi)(y_{p})$
$=K_{p}Q_{p,n}(\varphi)(y_{p}).$
Therefore by induction we have $Q_{p,n}^{\ast}(\varphi)=K_{p}Q_{p,n}(\varphi)$ for $p\in[0\,{:}\,n-1]$ and $Q_{n,n}^{\ast}(\varphi)=Q_{n,n}(\varphi)$ by definition, as required. ∎
Proposition A.3 (Twisting kernels equivalence).
Let $R:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$ and $K:(\mathsf{Y},\mathcal{Z})\rightarrow[0,1]$ be two Markov kernels for measurable spaces $(\mathsf{Y},\mathcal{Y})$ and $(\mathsf{Z},\mathcal{Z})$, and let $H$ be a non-negative real-valued function. If $R^{K(H)}$ and $K^{H}$ are twisted Markov kernels then $R^{K(H)}K^{H}=(RK)^{H}$.
@@
Note that since $K(H)$ is $L^{1}$-integrable w.r.t. $R(x,\cdot)$ by definition and $R\{K(H)\}=RK(H)$ then we have $H$ is $L^{1}$-integrable w.r.t. $RK(x,\cdot)$ for $x\in\mathsf{X}$.
@@
$R^{K(H)}K^{H}(x,\mathrm{d}z)=\begin{cases}\int_{\mathsf{Y}}\frac{R(x,\mathrm{d}y)K(H)(y)}{RK(H)(x)}K^{H}(y,\mathrm{d}z),&\text{if}~RK(H)(x)>0\\ QK^{H}(x,\mathrm{d}z),&\text{otherwise}.\end{cases}$
@@
by Proposition A.1. As for the second case, $QK^{H}$ is just another arbitrary Markov kernel, and hence $R^{K(H)}K^{H}=(RK)^{H}$. ∎
Proposition A.4 (Simplification of two $t$-knots).
@@
Let $\mathcal{M}=(M_{0:n},G_{0:n})$. By the knot definition, the model $\mathcal{K}_{2}\ast\mathcal{K}_{1}\ast\mathcal{M}=(M_{0:n}^{(2)},G_{0:n}^{(2)})$ has
@@
and the remaining kernels and potentials are the same as those in $\mathcal{M}$. By Proposition A.3 we can state $M_{t+1}^{(2)}=(K_{2}K_{1})^{G_{t}}M_{t+1}$. Therefore the kernels and potential of $\mathcal{K}_{2}\ast\mathcal{K}_{1}\ast\mathcal{M}$ are equal to that of $\mathcal{K}\ast\mathcal{M}$ where $\mathcal{K}=(t,R_{2},K_{2}K_{1})$. ∎
Proposition A.5 (Completion of a knotset).
@@
By Proposition A.3 we can state $V_{0}^{K_{0}(G_{0})}K_{0}^{G_{0}}=(V_{0}K_{0})^{G_{0}}$ and $R_{p}^{K_{p}(G_{p})}K_{p}^{G_{p}}=(R_{p}K_{p})^{G_{p}}$ for $p\in[1\,{:}\,n-1]$. Hence, by compatibility
@@
Proposition A.6 (Adapted knotset equivalence).
@@
Let $\mathcal{M}_{1}=\mathcal{K}^{\diamond}_{\ast}\ast\mathcal{M}^{\ast}=\mathcal{K}^{\diamond}_{\ast}\ast\mathcal{K}\ast\mathcal{M}$. The model $\mathcal{M}^{\ast}=(M_{0:n}^{\ast},G_{0:n}^{\ast})$ follows Proposition 3.9 and since $\mathcal{K}^{\diamond}_{\ast}$ is the adapted knotset for $\mathcal{M}^{\ast}$, we can state $\mathcal{M}_{1}=(M_{1,0:n},G_{1,0:n})$ where the kernels are $M_{1,0}=\delta_{0}$, $M_{1,p}=(M_{p-1}^{\ast})^{G_{p-1}^{\ast}}$ for $p\in[1\,{:}\,n-1]$, and $M_{1,n}=(M_{n-1}^{\ast})^{G_{n-1}^{\ast}}M_{n}$, whilst the potentials are $G_{1,p}=M_{p}^{\ast}(G_{p}^{\ast})$ for $p\in[0\,{:}\,n-1]$ and $G_{1,n}=G_{n}$. Noting that $R_{p}K_{p}=M_{p}$, $(M_{0}^{\ast})^{G_{0}^{\ast}}=R_{0}^{K_{0}(G_{0})}$, and $(M_{p}^{\ast})^{G_{p}^{\ast}}=(K_{p-1}^{G_{p-1}}R_{p})^{K_{p}(G_{p})}=(K_{p-1}^{G_{p-1}})^{M_{p}(G_{p})}R_{p}^{K_{p}(G_{p})}$ for $p\in[1\,{:}\,n-1]$ by Proposition A.3 we can state

### v2

@@ Appendix A Proof of main results
Appendix B Supporting results
Proposition B.1 (Kernel untwisting).
@@ Appendix B Supporting results
$K(H)\cdot K^{H}(\varphi)=K(H\cdot\varphi),$
$\lambda$-a.e. for any measure $\lambda$ on $(\mathsf{X},\mathcal{X})$.
@@
Second case. If $x\in\mathsf{X}_{0}$ we make three observations. Firstly, $K(H)(x)K^{H}(\varphi)(x)=K(H)(x)K(\varphi)(x)$ by definition. Secondly, $H=0$ almost surely w.r.t. $K(x,\cdot)$ since $H$ is non-negative. Hence, $K(H\cdot\varphi)(x)=0$ using the standard measure-theoretic convention $0\times\infty=0$ (see for example, Billingsley, 1995, p. 199). Lastly, $K(H)(x)K(\varphi)(x)=0$ for $K(\varphi)(x)\in[-\infty,\infty]$ by the same convention, since $K(H)(x)=0$. Therefore, $K(H)(x)K^{H}(\varphi)(x)=K(H\cdot\varphi)(x)=0$ for almost every $x\in\mathsf{X}_{0}$. ∎
Proposition B.2 (Form of $Q_{p,n}$ with knots).
Suppose $\mathcal{M}^{\ast}=\mathcal{K}\ast\mathcal{M}$ with $Q_{p,n}^{\ast}$ terms for $p\in[0\,{:}\,n]$. For any measure $\lambda_{p}$ on $(\mathsf{Y}_{p},\mathcal{Y}_{p})$, if $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$ and $\varphi:\mathsf{X}_{n}\rightarrow[-\infty,\infty]$ then $Q_{p,n}^{\ast}(\varphi)=K_{p}Q_{p,n}(\varphi)$ $\lambda_{p}$-a.e. for $p\in[0\,{:}\,n-1]$ and $Q_{n,n}^{\ast}(\varphi)=Q_{n,n}(\varphi)$.
@@
$Q_{p}^{\ast}(\varphi_{p})$ $=K_{p-1}(G_{p-1})\cdot K_{p-1}^{G_{p-1}}R_{p}(\varphi_{p})$ (10)
$=K_{p-1}\{G_{p-1}\cdot R_{p}(\varphi_{p})\},$
$\lambda_{p-1}$-a.e. by Proposition B.1. As for $Q_{n}^{\ast}$ we have
$Q_{n}^{\ast}(\varphi)$ $=K_{n-1}(G_{n-1})\cdot K_{n-1}^{G_{n-1}}M_{n}(\varphi)$
$=K_{n-1}\{G_{n-1}\cdot M_{n}(\varphi)\}$
$=K_{n-1}Q_{n}(\varphi),$
$\lambda_{n-1}$-a.e. by Proposition B.1. Similarly, it follows that
@@
Now, assume $Q_{p+1,n}^{\ast}(\varphi)=K_{p+1}Q_{p+1,n}(\varphi)$ for a given $p\in[0\,{:}\,n-2]$. Then by (10)
$Q_{p,n}^{\ast}(\varphi)$ $=Q_{p+1}^{\ast}Q_{p+1,n}^{\ast}(\varphi)$
$=\int K_{p}(\cdot,\mathrm{d}x_{p})G_{p}(x_{p})R_{p+1}(x_{p},\mathrm{d}y_{p+1})K_{p+1}(y_{p+1},\mathrm{d}x_{p+1})Q_{p+1,n}(\varphi)(x_{p+1})$
$=\int K_{p}(\cdot,\mathrm{d}x_{p})G_{p}(x_{p})M_{p+1}(x_{p},\mathrm{d}x_{p+1})Q_{p+1,n}(\varphi)(x_{p+1})$
$=K_{p}Q_{p+1}Q_{p+1,n}(\varphi)$
$=K_{p}Q_{p,n}(\varphi),$
$\lambda_{p}$-a.e. Therefore by induction we have $Q_{p,n}^{\ast}(\varphi)=K_{p}Q_{p,n}(\varphi)$ $\lambda_{p}$-a.e. for $p\in[0\,{:}\,n-1]$, and $Q_{n,n}^{\ast}(\varphi)=Q_{n,n}(\varphi)$ since $Q_{n,n}^{\ast}=Q_{n,n}=\mathrm{Id}$ by definition. ∎
Proposition B.3 (Twisting kernels equivalence).
Let $R:(\mathsf{X},\mathcal{Y})\rightarrow[0,1]$ and $K:(\mathsf{Y},\mathcal{Z})\rightarrow[0,1]$ be two Markov kernels for measurable spaces $(\mathsf{Y},\mathcal{Y})$ and $(\mathsf{Z},\mathcal{Z})$, and let $H$ be a non-negative real-valued function. If $R^{K(H)}$ and $K^{H}$ are twisted Markov kernels then $R^{K(H)}K^{H}=(RK)^{H}$ $\lambda$-a.e. for any measure $\lambda$ on $(\mathsf{X},\mathcal{X})$.
@@
Note that since $K(H)$ is integrable w.r.t. $R(x,\cdot)$ by definition and $R\{K(H)\}=RK(H)$ then we have $H$ integrable w.r.t. $RK(x,\cdot)$ for $x\in\mathsf{X}$.
@@
$R^{K(H)}K^{H}(x,\mathrm{d}z)=\begin{cases}\int_{\mathsf{Y}}\frac{R(x,\mathrm{d}y)K(H)(y)}{RK(H)(x)}K^{H}(y,\mathrm{d}z)&\text{if}~RK(H)(x)>0,\\ RK^{H}(x,\mathrm{d}z),&\text{if}~RK(H)(x)=0.\end{cases}$
@@
by Proposition B.1. As for the second case, $RK(H)(x)=0$ implies $K(H)=0$ almost surely w.r.t. $R$, hence $RK^{H}=RK$, by definition of $K^{H}$. Thus, $R^{K(H)}K^{H}=(RK)^{H}$ $\lambda$-a.e. ∎
Proposition B.4 (Simplification of two $t$-knots).
@@
Let $\mathcal{M}=(M_{0:n},G_{0:n})$. By repeated use of Definition 3.3, the model $\mathcal{K}_{2}\ast\mathcal{K}_{1}\ast\mathcal{M}=(M_{0:n}^{(2)},G_{0:n}^{(2)})$ has
@@
and the remaining kernels and potentials are the same as those in $\mathcal{M}$. By Proposition B.3 we can state $M_{t+1}^{(2)}=(K_{2}K_{1})^{G_{t}}M_{t+1}$. Therefore the kernels and potential of $\mathcal{K}_{2}\ast\mathcal{K}_{1}\ast\mathcal{M}$ are equal to that of $\mathcal{K}\ast\mathcal{M}$ where $\mathcal{K}=(t,R_{2},K_{2}K_{1})$. ∎
Proposition B.5 (Completion of a knotset).
@@
By Proposition B.3 we can state $V_{0}^{K_{0}(G_{0})}K_{0}^{G_{0}}=(V_{0}K_{0})^{G_{0}}$ and $R_{p}^{K_{p}(G_{p})}K_{p}^{G_{p}}=(R_{p}K_{p})^{G_{p}}$ for $p\in[1\,{:}\,n-1]$. Hence, by compatibility
@@
Proposition B.6 (Adapted knotset equivalence).
@@
Consider $\mathcal{K}=(R_{0:n-1},K_{0:n-1})$, $\mathcal{M}=(M_{0:n},G_{0:n})$, and let $\mathcal{M}_{1}=\mathcal{K}^{\diamond}_{\ast}\ast\mathcal{M}^{\ast}=\mathcal{K}^{\diamond}_{\ast}\ast\mathcal{K}\ast\mathcal{M}$. The model $\mathcal{M}^{\ast}=(M_{0:n}^{\ast},G_{0:n}^{\ast})$ follows Proposition 3.9 and since $\mathcal{K}^{\diamond}_{\ast}$ is the adapted knotset for $\mathcal{M}^{\ast}$, we can state $\mathcal{M}_{1}=(M_{1,0:n},G_{1,0:n})$ where the kernels are $M_{1,0}=\delta_{0}$, $M_{1,p}=(M_{p-1}^{\ast})^{G_{p-1}^{\ast}}$ for $p\in[1\,{:}\,n-1]$, and $M_{1,n}=(M_{n-1}^{\ast})^{G_{n-1}^{\ast}}M_{n}$, whilst the potentials are $G_{1,p}=M_{p}^{\ast}(G_{p}^{\ast})$ for $p\in[0\,{:}\,n-1]$ and $G_{1,n}=G_{n}$. Noting that $R_{p}K_{p}=M_{p}$, $(M_{0}^{\ast})^{G_{0}^{\ast}}=R_{0}^{K_{0}(G_{0})}$, and $(M_{p}^{\ast})^{G_{p}^{\ast}}=(K_{p-1}^{G_{p-1}}R_{p})^{K_{p}(G_{p})}=(K_{p-1}^{G_{p-1}})^{M_{p}(G_{p})}R_{p}^{K_{p}(G_{p})}$ for $p\in[1\,{:}\,n-1]$ by Proposition B.3 we can state